OpenStudy (anonymous):

I have a few questions I need help with. Here is the first one --------------------- On the blueprint of a house, 24 millimeters represents 5 meters. The length of the living room is 30 millimeters on the blueprint. What is the actual length of the living room?

OpenStudy (janu16):

24/5 = 30/x Solve for x. 24x=150 x=6.25 m

OpenStudy (janu16):

IS THT ONE OF THE CHOCIES?

OpenStudy (janu16):

@rikkibracamonte1999